1.Determine the pH during the titration of 36.6 mL of 0.304 M ethylamine (C2H5NH2 , Kb = 4.3×10-4) by 0.304 M HI at the following points.

(a) Before the addition of any HI

(b) After the addition of 16.1 mL of HI

(c) At the titration midpoint

(d) At the equivalence point

(e) After adding 51.2 mL of HI

b.Determine the pH during the titration of 61.4 mL of 0.450 M nitrous acid (Ka = 4.5×10-4) by 0.450 M NaOH at the following points.

(a) Before the addition of any NaOH

(b) After the addition of 15.0 mL of NaOH

(c) At the half-equivalence point (the titration midpoint)

(d) At the equivalence point

(e) After the addition of 92.1 mL of NaOH
Answer #1

ats NH + Ho - {H_NH + + oH Il 0.304 -1 0 0 X I tx | tX 0.304-X 1 X [ats NH₃ + ] LOH] [&HENI, k = 4.34104 = 2 0-304-X For weak< HsNt molarity = 0.304 M HI: Molanity = 0.304M volume = 36-6m 120.0366L volume = 16.1m/= 0.01612 no ot moles=0.304xcoolel 0.CH₃NH₂ + weak base ItI - strong acid {H NHI salt o Jo.0111264 1-o-oll1 204 0.011264 -0.011264 to-0111284 0.0 111264 Total vol6 HNO: Molarity = 0.45m I tx & HND Ho - Het + nor I boles | 0 To 6.45* - x x [Hot] [wog] tvo] NaOH Molanity = 0.45M volumes 6☺ At mid pornt pH apka oka = 3-347 so pH = 3-347 @ at equivalence point moles of ANO = moles of 0.0 Nach { Nang} = = 0.02763pH formulas for salts weak acid+ strong base -salft ito pH = ₂ [pkw tpka + logc] strong acid tweakbase salf tho pH = ₃ [ pk